# Think, Act, and Ask: Open-World Interactive Personalized Robot Navigation

### v2

@@ II Related Work
Dialogue Agents for Robots. Dialogue agents enable human-machine conversations through natural language [12, 13, 14, 15]. Recently, LLMs have been widely used in robots [16]. InnerMonologue [17] used an LLM to form an inner monologue style to ask questions. PromptCraft [18] explored prompt engineering practices for robot dialogues with ChatGPT. KNOWNO[19] measured the uncertainty of LLM planners for agents to ask for help when needed. In contrast, we use LLMs to operate different robot modules and frame it as a sequential decision-making problem.
Leverage Language Feedback. Autonomous agents that know how to adapt to language feedback can make more reliable decisions. Some research emphasizes real-time robot plan adjustments [20, 21]. Others harness language hints for assistance [22, 4] and human-machine collaboration [23, 24]. However, no study has comprehensively compared different feedback types in the same navigation context.
Personalized Human-Robot Interaction. Building personalized robots is an active research area in human-robot interaction [25, 26, 27]. There are many works focused on enhancing personalized experience through non-verbal communications [28, 29, 30] and better interactive service design [31]. Personalized dialogue systems also gained increasing interest [32, 33, 34, 35], where the persona is taken as conditional input to produce more characterized and sociable conversations. To our knowledge, previous work has not investigated interactive personalized navigation tasks.
@@ III-A The ZIPON Task
ZIPON is a generalized type of zero-shot object navigation [4]. Let denote the set of all test scenes, and denote the set of all personalized goals. Each goal contains a tuple , where type is the class label for the object (e.g., ‘bed’, ‘chair’), name is the personalized expression (e.g. ‘Alice’s bed’, ‘chair bought from Amazon’) which is unique for every goal, room is the name of the room (e.g., ‘Alice bedroom’, ‘living room’) where is located in, and FB is a dictionary to store all types of user feedback information (See Sec. IV-A). A navigation episode is a tuple , where , , is the initial pose of the agent for current . The input observations are RGB-D images. Starting from , the agent needs to find by taking a sequence of primitive actions, where the action space is ={TurnLeft 15, TurnRight 15, MoveForward 0.25m, Talk}. The first three are navigation actions, and the last one is a communicative action with the dialogue content to be generated by the agent. During the evaluation for , the agent can either move in the environment with navigation actions or interact with the user with Talk for more information. Whenever the agent believes it has reached the goal, it must issue Talk to stop moving and confirm with the user. The is terminated when the robot successfully finds the or a maximum number of interaction attempts (i.e., the total number of Talk actions) is attained. If the agent is within meters of and meets visibility criteria, the is successful.
@@ III-B Proposed ORION Framework
We propose ORION to solve ZIPON. As shown in Fig. 2, this framework comprises six modules: a control module to perform navigation movements, a semantic map module for natural language indexing, an open-vocabulary detection module to detect any objects with language descriptions, an exploration module to search the room, a memory module to store crucial information from user feedback, and an interaction module to talk with users. Central to these, the Large Language Model (LLM) serves as the primary controller, making sequential actions based on its strong reasoning abilities.
@@ III-B2 Operated Modules
Control Module. It contains two low-level navigation actions, Move(num) and Rotate(num), that cover the primitive navigation actions in . Concretely, Move(1) is MoveForward 0.25m, Rotate(1) denotes TurnRight 15 and TurnLeft 15, respectively. In addition, it has two high-level actions: goto_point(point), that sequences low-level navigation actions to navigate to a point on a 2D occupancy map, and goto_object(obj_id) that directly navigates to an object with the given id.
@@
Exploration Module. This module uses frontier-based exploration (FBE) [39] to help the agent search unseen scenes. Following [3], the agent starts with a 360 spin and builds a 2D occupancy grid map for exploration. The LLM action for this module is search_object(obj_str), where FBE determines the next frontier points for the control module to take goto_point and the detection module to perceive on-the-fly with detect_object. Upon detecting objects, this module returns the detection message. If invoked again, it continues to explore the room until FBE stops.
@@ IV-B3 Evaluation Results
Ablations and more analysis. Tab. II’s lower section presents ablation results for ORION in the mixed feedback setting. Here, ‘w/o mem’ excludes the additional memory feature; ‘w/o exp’ omits the exploration module but retains the 360 spin; ‘w/o det’ replaces the grounded-SAM model with a basic k-patch mechanism [3] where the CLIP matches texts to image patches; and ‘w/o map’ removes the semantic map module. The results emphasize the crucial role of the semantic map, marked by the most significant SR drop in ORION w/o map. Both ORION w/o exp and ORION w/o det yield comparable outcomes, suggesting the importance of a robust detection model paired with active exploration. Interestingly, the memory module has a limited impact, even with its capability to retain user information. We hypothesize this is because, during the zero-shot evaluation, previously stored goals aren’t retested. To further explore this, we conduct a second time ZIPON evaluation using the memory accumulated from the first time. Consequently, ORION gets scores of 91.5% in SR, 63.9% in SPL, and 65.3% in SIT.
@@ IV-C1 Experimental Setup
| Corrective | 67.5 | 51.8 | 53.6 |
| Descriptive | 67.5 | 48.6 | 48.0 |
| Landmark | 72.5 | 66.8 | 54.8 |
| Procedural | 72.5 | 62.8 | 61.9 |

### v4

@@ II Related Work
Dialogue Agents for Robots. Dialogue agents enable human-machine conversations through natural language [12, 13, 14, 15, 16]. A large number of earlier works have studied language use in human-robot dialogue [17, 18, 19]. Recently, LLMs have been widely used in robots [20]. InnerMonologue [21] used an LLM to form an inner monologue style to ask questions. PromptCraft [22] explored prompt engineering practices for robot dialogues with ChatGPT. KNOWNO[23] measured the uncertainty of LLM planners for agents to ask for help when needed. In contrast, we use LLMs to operate different robot modules and frame it as a sequential decision-making problem.
Leverage Language Feedback. In human-robot dialogue, robots that can learn from and adapt to language feedback can make more reliable decisions [24, 25, 26]. Previous works have emphasized real-time robot plan adjustments [27, 28]. Others harness language instructions for assistance [29, 4], task learning [30], and human-machine collaboration [31, 32, 33]. However, no study has comprehensively compared different feedback types in the navigation context.
Personalized Human-Robot Interaction. Building personalized robots is an active research area in human-robot interaction [34, 35, 36]. There are many works focused on enhancing personalized experience through non-verbal communications [37, 38, 39] and better interactive service design [40, 41]. Personalized dialogue systems also gained increasing interest [42, 43, 44, 45], where the persona is taken as conditional input to produce more characterized and sociable conversations. To our knowledge, previous work has not investigated interactive personalized navigation tasks.
@@ III-A The ZIPON Task
ZIPON is a generalized type of zero-shot object navigation [4]. Let denote the set of all test scenes, and denote the set of all personalized goals. Each goal contains a tuple , where type is the class label for the object (e.g., ‘bed’, ‘chair’), name is the personalized expression (e.g. ‘Alice’s bed’, ‘chair bought from Amazon’) which is unique for every goal, room is the name of the room (e.g., ‘Alice bedroom’, ‘living room’) where is located in, and FB is a dictionary to store all types of user feedback information (See Sec. IV-A). A navigation episode is a tuple , where , , is the initial pose of the agent for current . The input observations are RGB-D images. Starting from , the agent needs to find by taking a sequence of primitive actions, where the action space is ={TurnLeft 15∘, TurnRight 15∘, MoveForward 0.25m, Talk}. The first three are navigation actions, and the last one is a communicative action with the dialogue content to be generated by the agent. During the evaluation for , the agent can either move in the environment with navigation actions or interact with the user with Talk for more information. Whenever the agent believes it has reached the goal, it must issue Talk to stop moving and confirm with the user. The is terminated when the robot successfully finds the or a maximum number of interaction attempts (i.e., the total number of Talk actions) is attained. If the agent is within meters of and meets visibility criteria, the is successful.
@@ III-B Proposed ORION Framework
We propose ORION to solve ZIPON. As shown in Fig. 2, this framework comprises six modules: a control module to perform navigation movements, a semantic map module for natural language indexing, an open-vocabulary detection module to detect any objects with language descriptions, an exploration module to search the room, a memory module to store crucial information from user feedback, and an interaction module to talk. Central to these, the Large Language Model (LLM) serves as the primary controller, making sequential actions based on its strong reasoning abilities.
@@ III-B2 Operated Modules
Control Module. It contains two low-level navigation actions, Move(num) and Rotate(num), that cover the primitive navigation actions in . Concretely, Move(1) is MoveForward 0.25m, Rotate(1) denotes TurnRight 15∘ and TurnLeft 15∘, respectively. In addition, it has two high-level actions: goto_point(point), that sequences low-level navigation actions to navigate to a point on a 2D occupancy map, and goto_object(obj_id) that directly navigates to an object with the given id.
@@
Exploration Module. This module uses frontier-based exploration (FBE) [49] to help the agent search unseen scenes. Following [3], the agent starts with a 360∘ spin and builds a 2D occupancy grid map for exploration. The LLM action for this module is search_object(obj_str), where FBE determines the next frontier points for the control module to take goto_point and the detection module to perceive on-the-fly with detect_object. Upon detecting objects, this module returns the detection message. If invoked again, it continues to explore the room until FBE stops.
@@ IV-B3 Evaluation Results
Ablations and more analysis. Tab. II’s lower section presents ablation results for ORION in the mixed feedback setting. Here, ‘w/o mem’ excludes the additional memory feature; ‘w/o exp’ omits the exploration module but retains the 360∘ spin; ‘w/o det’ replaces the grounded-SAM model with a basic k-patch mechanism [3] where the CLIP matches texts to image patches; and ‘w/o map’ removes the semantic map module. The results emphasize the crucial role of the semantic map, marked by the most significant SR drop in ORION w/o map. Both ORION w/o exp and ORION w/o det yield comparable outcomes, suggesting the importance of a robust detection model paired with active exploration. Interestingly, the memory module has a limited impact, even with its capability to retain user information. We hypothesize this is because, during the zero-shot evaluation, previously stored goals aren’t retested. To further explore this, we conduct a second time ZIPON evaluation using the memory accumulated from the first time. Consequently, ORION gets scores of 91.5% in SR, 63.9% in SPL, and 65.3% in SIT.
@@ IV-C1 Experimental Setup
| Corrective | 67.5∗ | 51.8∗ | 53.6∗ |
| Descriptive | 67.5∗ | 48.6 | 48.0 |
| Landmark | 72.5∗ | 66.8∗ | 54.8∗ |
| Procedural | 72.5∗ | 62.8∗ | 61.9∗ |